OpenStudy (anonymous):

Sam has a total of 58 DVDs and CDs. If the number of CDs is two more than three times the number of DVDs, how many CDs does he have?

OpenStudy (anonymous):

Help anyone?

OpenStudy (anonymous):

Let D be the number of DVDs. Then C, the number of CDs is C = 3D + 2 The total is 58, so C + D = 58 3D + 2 + D = 58 4D = 56 D = 14 C = 3(14) + 2 = 42 + 2 = 44 So he has 14 DVDs and 44 CDs

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

welcome ;)